OpenStudy (anonymous):

what 5 × ( + 4) = 30

OpenStudy (alexandervonhumboldt2):

does the question say 5*(x+4)=30?

iYuko (iyuko):

Insufficient information to solve.

OpenStudy (anonymous):

5 x ( + 4) = 30 AlexandervonHumboldt2

iYuko (iyuko):

If it says: 5*(x+4)=30 Divide each side by 5 x+4=6 subtract 4 from each side x=2